A 900kg car is travelling at 30km/hour in the positive x-direction when the driver fails to notice that the vehicle in front of then has stopped - they drive into the back of the stopped car. Both cars are stationary after the collision. What was the change on momentum of the first car?



1
Expert's answer
2021-05-05T10:53:47-0400

"p_1=-mv,~p_2=0,"

"\Delta p=p_2-p_1=0-(-mv)=mv,"

"\Delta p=900\cdot \frac{30}{3.6}=7500~\frac{kg\cdot m}{s}."